OpenStudy (anonymous):

I need math help! I cant figure out how to do it! here is the question: The total of 5.68 times a number and 132.7 is the number minus 29.228. Find the number.

OpenStudy (anonymous):

need some help?

Directrix (directrix):

Let x be the number .568*x + 132.7 = x - 29.228 @sunshinedoo4u Do you think that this equation is correct?

OpenStudy (anonymous):

5.68x+132.7=x-29.228 i dont even know where to start.

Directrix (directrix):

5.68x+132.7=x-29.228 Subtract 1x from both sides.

Directrix (directrix):

5.68x+132.7= x-29.228 -x -x ---------------------------- 4.68x + 132.7 = - 29.228

Directrix (directrix):

Next up: Subtract 132.7 from both sides.